A battery of 10 volt carries 20,000 C of charge through a resistance of 20 Ω. The work done in 10 seconds is
(a) 2 × 103 joule
(b) 2 × 105joule
(c) 2 × 104 joule
(d) 2 × 102 joule

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: b Explanation: (b) W= qV= 20000 × 10 = 2,00, 000 = 2 × 105 J